OpenStudy (anonymous):

Solve by completing the square: x(squared) + 2x -15 And 6 + 4x - x(squared)

Directrix (directrix):

Steps for completing the square here; http://www.regentsprep.org/regents/math/algtrig/ate12/completesqlesson.htm

Directrix (directrix):

x^2 + 2x + ___ = 15 + _____ x^2 + 2x + __ 1 __ = 15 + __1__ (x + 1) ^2 = 16 Take square roots (x + 1 ) = 4 or x + 1 = -4 x = 3 or x = -5

Directrix (directrix):

You do this one: 6 + 4x - x^2

OpenStudy (anonymous):

@Sarkiny

OpenStudy (anonymous):

Thank you! x^2 - 4x - 6 (x - 2)^2 - 4 - 6 (x - 2)^2 - 10 I think this is the solution only on one side because in this question they don't ask us to make a second side. So the first that you did is gonna be (x + 1)^2 -16 right?

OpenStudy (anonymous):

yes,it should be (x+1)^2-16 if the solution is only one side @Sarkiny